# Moments and the Basic Equilibrium Equations

### v1

@@ 2. Moments of a Weighted System
Part (1) follows directly from lemma 2.2 and the definition of $G$. For part (2), note that $\mu_{1}$ is constant according to lemma 2.2, so we have that $\mu_{1}(p)=\mu_{1}(x)$. But the total weight of the weighted system $\left(X\setminus\{x\},w\big{|}_{X\setminus\{x\}}\right)$ is $-w(x)\neq 0$ and its first moment at the point $x$ is $\mu_{1}(x)$. So, by applying part (1) to the latter system with $p=x$, we obtain part (2). ∎
@@ 3. Equilibrium Equations
(3.1) $\quad\overrightarrow{F}_{x,y}=\varphi_{x,y}\overrightarrow{xy},\quad\text{with}\ \varphi_{x,y}:\prod_{x\in X}\mathbb{A}\longrightarrow K\ \text{arbitrary}.$
Notice that the expression (3.1) is commonly adopted to describe internal forces in Classical Mechanics, together with conditions on the $\varphi_{x,y}$ in order to comply with hypotheses on homogeneity and isotropy of space [6]. For our purposes, though, it is necessary neither to make these additional assumptions, nor to impose the validity of Newton’s third law of motion, which corresponds to equalities $\varphi_{x,y}=\varphi_{y,x}$ for all $x,y\in X$.
@@ 4. The space of weighted systems with $\mu_{1}={\overrightarrow{O}}$
From now on we let $X=\{x_{1},\ldots,x_{n}\}$. Denote the dimension of the affine closure of $X$ by $\text{dim}(X)$, or just $d$, and call it the dimension of $X$. The difference between the maximum possible value for $\text{dim}(X)$, which is $n-1$, and $d$, that is $c=(n-1)-d$, is called the codimension of $X$, and we write $\text{codim}(X)=c$. A weight function $w:X\longrightarrow\mathbb{R}$ may alternatively be represented by the vector $W\in\mathbb{R}^{n}$ whose coordinates are $w(x_{j})=w_{j}$.
@@
that is, $d+1$. The nullity of $\mathcal{X}$ is thus $n-(d+1)=(n-1)-d=c$. This is the dimension of $\mathbb{W}_{0}(X)$. ∎
@@
As we shall soon see, a key role is played by configurations whose codimension is 1. Borrowing the terminology from Celestial Mechanics, such configurations will be designated as Dziobek configurations, calling them Dziobek subconfigurations when they are proper subsets of a larger configuration.
Notice that a configuration $X$ has dimension $d$ if and only if $X$ contains an ascending chain of simplexes $S_{0}\subset\ldots\subset S_{d-1}\subset S_{d}$, where each $S_{k}$ is $k$-dimensional, $k=0,\ldots,d-1,d$. We refer to any simplex of dimension $d$ contained in $X$ as a core of $X$. There are $n-(d+1)=c$ points in $X\setminus S_{d}$, so we have exactly $c$ distinct Dziobek subconfigurations $D^{(1)},\ldots,D^{(c)}$ containing the core $S_{d}$. Figure 1 depicts one such arrangement of simplexes and Dziobek subconfigurations of $X$. We call each such diagram a Dziobek tree for $X$.
@@
which is defined by placing zeros at the coordinates corresponding to the $n-(d+2)=c-1$ points in $X\setminus D^{(l)}$.
@@
We must have that $w_{d+l+1}\neq 0$, for all $l=1,\ldots,c$, otherwise the above equations would provide a nontrivial linear combination of the vectors $\overrightarrow{x_{1}x_{j}}$, $j=2,\ldots,d+1$, contradicting the hypothesis that $S_{d}$ is a core. By adding $c-1$ zeros as previously explained, we embed each of the spanning vectors for the Dziobek subconfigurations $D^{(l)}$ into $\mathbb{W}_{0}(X)$, thus obtaining the collection of $c$ weight vectors
@@
Considering all possible Dziobek trees for $X$, we see from proposition 4.3 that the maximum number of bases obtainable by our method is (${n\choose d+1}={n\choose c}$. The maximum is attained when each subconfiguration of $X$ formed by $d+1$ points is a core of $X$. In this case, we call $X$ a generic configuration. A configuration is termed nongeneric if it is not generic. The type of a nongeneric configuration refers to the number of its subconfigurations with $d+1$ points that are not cores. We illustrate the possibilities for five points on a plane in figure 2. An application is made in example 5.3.
Figure 2. Planar configurations of five points: generic (left), nongeneric of type $1$ (left center), nongeneric of type 2 (center right), and nongeneric of type $4$ (right).
@@
Let $X=\{x_{1},\ldots,x_{n}\}$, and let $X_{j}$ be the subconfiguration $X\setminus\{x_{j}\}$. Denote by $\mathcal{X}_{j}$ the configuration matrix of $X_{j}$. $X$ has codimension $1$ if and only if $\mathbb{W}_{0}(X)$ is spanned by the weight vector
$W=(-\Delta_{1},\Delta_{2},-\Delta_{3},\ldots,(-1)^{n}\Delta_{n})$
@@
The (oriented) volume of a configuration $X$ of codimension $c$ is the determinant $\det(\mathcal{X})$ of its configuration matrix $\mathcal{X}$ if $c=0$, and zero otherwise. Let $X_{j_{1}\ldots j_{\beta}}$ denote the subconfiguration $X\setminus\{x_{j_{1}},\ldots,x_{j_{\beta}}\}$, where $j_{1}<\ldots<j_{\beta}$ and let $\mathcal{X}_{j_{1}\ldots j_{\beta}}$ be the configuration matrix of $X_{j_{1}\ldots j_{\beta}}$. Denote the volume of $X_{j_{1}\ldots j_{\beta}}$ by $\Delta_{j_{1}\ldots j_{\beta}}$. The construction of bases for $\mathbb{W}_{0}(X)$ using the volume barycentric coordinates $\Delta_{j_{1}\ldots j_{\beta}}$ is similar to the one done in the proof of proposition 4.3. For the sake of concreteness, we provide an example.
@@
Suppose $X_{4567}=\{x_{1},x_{2},x_{3}\}$ is a core of $X$, thus the corresponding volume, $\Delta_{4567}$, is nonzero. $X_{4567}$ is contained in four Dziobek subconfigurations: $X_{567}$, $X_{467}$, $X_{457}$, $X_{456}$. The Dziobek subconfigurations produce the respective weight vectors
@@
We end this section with the statement of a criterion for membership in $\mathbb{W}_{0}(X)$. The proof is followed by related remarks.
@@
Each of the above vectors corresponds to a Dziobek subconfiguration $\{x_{1},\ldots,x_{d+1},x_{l}\}$, with $l=d+2,\ldots,n$ ; we denote them by $W_{l}$, respectively.
@@ 5. Some Applications to central configurations
We call a configuration of particles (point masses, bodies) in $\mathbb{R}^{N}$ any weighted system $(X,m)$ where $m$ is a nonvanishing function which we traditionally refer to as mass. Recall the notation used in sections 3 and 4. The points of $X$ are denoted by $x_{i}$, $m_{i}$ is the value of $m$ at $x_{i}$, $i=1,\ldots,n$, the dimension of $X$ is $d$ and the codimension of $X$ is $c=(n-1)-d$. We may interpret $m$ as the weight vector $(m_{1},\ldots,m_{n})\in\mathbb{W}=\mathbb{R}^{n}$. Assuming the particles interact under some homogeneous potential, the corresponding acceleration vectors have the form
@@
Albouy [4] has shown that, if $a=0$, then any configuration is central, and if $a\neq 0$ and $d=n-1$, i.e, if $c=0$, then $X$ is a regular simplex for all values of $\mu_{0}$. In the cases $\mu_{0}\neq 0$, this assertion follows from lemma 4.1 and theorem 5.2 below (as in [4], we tacitly assume that $\lambda/\mu_{0}>0$).
@@
Next, let us consider the case $\mu_{0}\neq 0$. As proved in [4], $(X,m)$ is a central configuration if and only if $\overrightarrow{\gamma}_{O}$ is the null vector and $x_{O}=\text{bar}(X,m)$ is the center of mass of $(X,m)$. The central configuration equations can be put in the form
@@
The corollary follows immediately from theorems 4.25.2 and the fact that $m$ is nonvanishing. In the Dziobek case, $c=1$, the last statement in the corollary implies that $S_{ij}S_{kl}-S_{il}S_{jk}=0$ for any four indices $i,j,k,l$. When the indices are distinct, these identities are the well known equations (18) in [28]. By considering some equal indices in the case $n=4$ and $c=1$, for instance, we compute $S_{12}S_{24}-S_{14}S_{22}=0$, which gives
@@
    where the $+$ (resp. --) sign occurs if $x_{4}$ and $x_{5}$ are on opposite sides (resp. the same side) of $x_{1}x_{2}$.
@@
    In particular, for $a<-1/2$ and $m_{1}=m_{3}$, particle $x_{2}$ is at the midpoint of $x_{1}$ and $x_{3}$. Moreover, if the lines $x_{1}x_{2}$ and $x_{4}x_{5}$ are parallel, then, for all $m_{2}$, the configuration is an isosceles trapezoid and $m_{4}=m_{5}$.
@@
Item (a) follows from the first determinant in equations (5.5) and from $\frac{\Delta_{j4}}{\Delta_{j5}}=\pm\frac{h_{5}}{h_{4}}$. For (b), notice $h_{4}=h_{5}$ forces $x_{1}x_{2}$ to intersect $x_{4}x_{5}$ at the midpoint of $x_{4}$ and $x_{5}$. Apply the Perpendicular Bisector Theorem [4] and item (a) to draw the stated conclusions. To obtain the equation in (c), we substitute $S_{24}=0$ in the second determinant of (5.5) with $i=1,j=2,k=3,l=4$ and use the equation $\frac{\Delta_{14}}{\Delta_{34}}=\frac{s_{23}^{1/2}}{s_{12}^{1/2}}$. The remaining statements in (c) follow from the fact that $f(x)=x^{2a+1}-(\lambda/\mu_{0})x$, $x>0$, is an injective function for any $a<-1/2$, and from item (a). Indeed, $h_{4}=h_{5}$ by hypothesis, so $X$ is an isosceles trapezoid and, by symmetry, $S_{14}=S_{35}$ and $S_{34}=S_{15}$. Thus, by (a),
@@
References [12, 18] describe a family of planar central configurations of five bodies under the Newtonian gravitational potential (i.e., for $a=-3/2$) which fulfills the conditions of item (c) of proposition 5.4. In both [12] and [18], the masses are positive and the symmetry of $(X,m)$ is assumed from start. On the other hand, the hypothesis $s_{24}^{a}=\lambda/\mu_{0}$ is part of the conclusions in [12], which also discusses the changes in the shape of the isosceles trapezoid as $m_{2}$ changes. We note that reference [18] examines configurations considered in item (b) of proposition 5.4, as well as some symmetrical configurations with $m_{4}\neq m_{5}$ (none of which turns out to be central).
@@ 6. Generalized Dziobek Equations
The number of parameters defining $S$ is (${c+1\choose 2}$, which, according to formula (A.2.4) in appendix A, is precisely the number of constraints that must be imposed on the mutual distances of a configuration of $n$ points to guarantee it has a given codimension $c>0$. This coincidence is not fortuitous since, by theorem A.2.2, the vanishing of (${c+1\choose 2}$ Cayley-Menger determinants provide the necessary constraints. Dziobek’s approach via critical point theory then introduces (${c+1\choose 2}$ Lagrange multipliers. Dziobek also verified that the derivative of a Cayley-Menger determinant with respect to a squared mutual distance can be expressed simply as a constant times the product of a pair of oriented volumes of suitable subconfigurations [16, 21]. Aside from determining the exact value of the constant, this fact can be readily deduced from theorem 6.1. Our proof of theorem 6.1 was inspired by the elegant proof of proposition 8 in [4].

### v2

@@ 2. Moments of a Weighted System
Part (1) follows directly from lemma 2.2 and the definition of $G$. For part (2), note that $\mu_{1}$ is constant according to lemma 2.2, so we have that $\mu_{1}(p)=\mu_{1}(x)$. But the total weight of the weighted system $\left(X\setminus\{x\},w\big{|}_{X\setminus\{x\}}\right)$ is $w(x)\neq 0$ and its first moment at the point $x$ is $\mu_{1}(x)$. So, by applying part (1) to the latter system with $p=x$, we obtain part (2). ∎
@@ 3. Equilibrium Equations
(3.1) $\quad\overrightarrow{F}_{x,y}=\varphi_{x,y}\overrightarrow{xy},\quad\text{with }\ \varphi_{x,y}:\prod_{z\in X}\mathbb{A}\longrightarrow K\ \text{arbitrary}.$
Notice that the expression (3.1) is commonly adopted to describe internal forces in Classical Mechanics, together with conditions on the $\varphi_{x,y}$ in order to comply the homogeneity and isotropy of space [6]. For our purposes, though, it is necessary neither to make these additional assumptions, nor to impose the validity of Newton’s third law of motion, which corresponds to equalities $\varphi_{x,y}=\varphi_{y,x}$ for all $x,y\in X$.
@@ 4. The space of weighted systems with $\mu_{1}={\overrightarrow{O}}$
From now on we let $X=\{x_{1},\ldots,x_{n}\}$. Denote the dimension of the affine closure of $X$ by $\text{dim}(X)$, or just $d$, and call it the dimension of $X$. The difference between the maximum possible value for $\text{dim}(X)$, which is n1n-1, and $d$, that is c=(n1)dc=(n-1)-d, is called the codimension of $X$, and we write $\text{codim}(X)=c$. A weight function $w:X\longrightarrow\mathbb{R}$ may alternatively be represented by the vector $W\in\mathbb{R}^{n}$ whose coordinates are $w(x_{j})=w_{j}$.
@@
that is, $d+1$. The nullity of $\mathcal{X}$ is thus n(d+1)=(n1)d=cn-(d+1)=(n-1)-d=c. This is the dimension of $\mathbb{W}_{0}(X)$. ∎
@@
As we shall soon see, a key role is played by configurations whose codimension is 1. Borrowing the terminology from Celestial Mechanics, such a configuration will be designated as Dziobek configuration, or Dziobek subconfiguration, when it is a proper subset of a larger configuration.
Notice that a configuration $X$ has dimension $d$ if and only if $X$ contains an ascending chain of simplexes $S_{0}\subset\ldots\subset S_{d-1}\subset S_{d}$, where each $S_{k}$ is $k$-dimensional, $k=0,\ldots,d-1,d$. We refer to any simplex of dimension $d$ contained in $X$ as a core of $X$. There are n(d+1)=cn-(d+1)=c points in $X\setminus S_{d}$, so we have exactly $c$ distinct Dziobek subconfigurations $D^{(1)},\ldots,D^{(c)}$ containing the core $S_{d}$. Figure 1 depicts one such arrangement of simplexes and Dziobek subconfigurations of $X$. We call each such diagram a Dziobek tree for $X$.
@@
which is defined by placing zeros at the coordinates corresponding to the n(d+2)=c1n-(d+2)=c-1 points in $X\setminus D^{(l)}$.
@@
We must have that $w_{d+l+1}\neq 0$, for all $l=1,\ldots,c$, otherwise the above equations would provide a nontrivial linear combination of the vectors $\overrightarrow{x_{1}x_{j}}$, $j=2,\ldots,d+1$, contradicting the hypothesis that $S_{d}$ is a core. By adding c1c-1 zeros as previously explained, we embed each of the spanning vectors for the Dziobek subconfigurations $D^{(l)}$ into $\mathbb{W}_{0}(X)$, thus obtaining the collection of $c$ weight vectors
@@
Considering all possible Dziobek trees for $X$, we see from proposition 4.3 that the maximum number of bases obtainable by our method is ${n\choose d+1}={n\choose c}$. The maximum is attained when each subconfiguration of $X$ formed by $d+1$ points is a core of $X$. In this case, we call $X$ a generic configuration. A configuration is termed nongeneric if it is not generic. The type of a nongeneric configuration refers to the number of its subconfigurations with $d+1$ points that are not cores. We illustrate the possibilities for five points on a plane in figure 2. An application is made in example 5.3.
Figure 2. Planar configurations of five points: generic (left), nongeneric of type 11 (left center), nongeneric of type 2 (center right), and nongeneric of type 44 (right).
@@
Let $X=\{x_{1},\ldots,x_{n}\}$, and let $X_{j}$ be the subconfiguration $X\setminus\{x_{j}\}$. Denote by $\mathcal{X}_{j}$ the configuration matrix of $_{j}. XX$ has codimension $1$ if and only if $\mathbb{W}_{0}(X)$ is spanned by the weight vector
$W=(-\Delta_{1},\Delta_{2},-\Delta_{3},\ldots,(-1)^{n}\Delta_{n}),$
@@
The (oriented) volume of a configuration $X$ of codimension $c$ is the determinant $\det(\mathcal{X})$ of its configuration matrix $\mathcal{X}$ if $c=0$, and zero otherwise. Let $X_{j_{1}\ldots j_{\beta}}$ denote the subconfiguration $X\setminus\{x_{j_{1}},\ldots,x_{j_{\beta}}\}$, where $j_{1}<\ldots<j_{\beta}$, and let $\mathcal{X}_{j_{1}\ldots j_{\beta}}$ be the configuration matrix of $X_{j_{1}\ldots j_{\beta}}$. Denote the volume of $X_{j_{1}\ldots j_{\beta}}$ by $\Delta_{j_{1}\ldots j_{\beta}}$. The construction of bases for $\mathbb{W}_{0}(X)$ using the volume barycentric coordinates $\Delta_{j_{1}\ldots j_{\beta}}$ is similar to the one done in the proof of proposition 4.3. For the sake of concreteness, we provide an example.
@@
Suppose $X_{4567}=\{x_{1},x_{2},x_{3}\}$ is a core of $X$; the corresponding volume $\Delta_{4567}$ is, thus, nonzero. $X_{4567}$ is contained in four Dziobek subconfigurations: $X_{567}$, $X_{467}$, $X_{457}$, $X_{456}$. These Dziobek subconfigurations produce the respective weight vectors
@@
We end this section with a criterion for membership in $\mathbb{W}_{0}(X)$. The proof is followed by related remarks.
@@
Each of the above vectors corresponds to a Dziobek subconfiguration $\{x_{1},\ldots,x_{d+1},$ $x_{l}\}$, with $l=d+2,\ldots,n$ ; we denote the weight vector associated to $l$ by $W_{l}$.
@@ 5. Some Applications to central configurations
We call a configuration of particles (point masses, bodies) in $\mathbb{R}^{N}$ any weighted system $(X,m)$ where $m$ is a nonvanishing function which we traditionally refer to as mass. Recall the notation used in sections 3 and 4. The points of $X$ are denoted by $x_{i}$, $m_{i}$ is the value of $m$ at $x_{i}$, $i=1,\ldots,n$, the dimension of $X$ is $d$ and the codimension of $X$ is c=(n1)dc=(n-1)-d. We may interpret $m$ as the weight vector $(m_{1},\ldots,m_{n})\in\mathbb{W}=\mathbb{R}^{n}$. Assuming the particles interact under some homogeneous potential, the corresponding acceleration vectors have the form
@@
Albouy [4] has shown that, if $a=0$, then any configuration is central, and if $a\neq 0$ and d=n1d=n-1, i.e, if $c=0$, then $X$ is a regular simplex for all values of $\mu_{0}$. In the cases $\mu_{0}\neq 0$, this assertion follows from lemma 4.1 and theorem 5.2 below (as in [4], we tacitly assume that $\lambda/\mu_{0}>0$).
@@
Next, let us consider the case $\mu_{0}\neq 0$. As proved in [4], $(X,m)$ is a central configuration if and only if $\overrightarrow{\gamma}_{O}$ is the null vector and $x_{O}=\text{bar}(X,m)$ is the center of mass of $(X,m)$. The central configuration equations can thus be put in the form
@@
The previous corollary follows immediately from theorems 4.25.2 and the fact that $m$ is nonvanishing. In the Dziobek case, $c=1$, the last statement in the corollary implies that $S_{ij}S_{kl}-S_{il}S_{jk}=0$ for any four indices $i,j,k,l$. When the indices are distinct, these identities are the well known equations (18) in [28]. By considering some equal indices in the case $n=4$ and $c=1$, for instance, we compute $S_{12}S_{24}-S_{14}S_{22}=0$, which gives
@@
    where the ++ (resp. -) sign occurs if $x_{4}$ and $x_{5}$ are on opposite sides (resp. the same side) of $x_{1}x_{2}$.
@@
    In particular, for a<1/2a<-1/2 and $m_{1}=m_{3}$, particle $x_{2}$ is at the midpoint of $x_{1}$ and $x_{3}$. Moreover, if the lines $x_{1}x_{2}$ and $x_{4}x_{5}$ are parallel, then, for all $m_{2}$, the configuration is an isosceles trapezoid and $m_{4}=m_{5}$.
@@
Item (a) follows from the first determinant in equations (5.5) and from $\frac{\Delta_{j4}}{\Delta_{j5}}=\pm\frac{h_{5}}{h_{4}}$. For (b), notice $h_{4}=h_{5}$ forces $x_{1}x_{2}$ to intersect $x_{4}x_{5}$ at the midpoint of $x_{4}$ and $x_{5}$. Apply the Perpendicular Bisector Theorem [4] and item (a) to draw the stated conclusions. To obtain the equation in (c), we substitute $S_{24}=0$ in the second determinant of (5.5) with $i=1,j=2,k=3,l=4$ and use the equation $\frac{\Delta_{14}}{\Delta_{34}}=\frac{s_{23}^{1/2}}{s_{12}^{1/2}}$. The remaining statements in (c) follow from the fact that $f(x)=x^{2a+1}-(\lambda/\mu_{0})x$, $x>0$, is an injective function for any a<1/2a<-1/2, and from item (a). Indeed, $h_{4}=h_{5}$ by hypothesis, so $X$ is an isosceles trapezoid and, by symmetry, $S_{14}=S_{35}$ and $S_{34}=S_{15}$. Thus, by (a),
@@
References [12, 18] describe a family of planar central configurations of five bodies under the Newtonian gravitational potential (i.e., for a=3/2a=-3/2) which fulfills the conditions of item (c) of proposition 5.4. In both [12] and [18], the masses are positive and the symmetry of $(X,m)$ is assumed from start. On the other hand, the hypothesis $s_{24}^{a}=\lambda/\mu_{0}$ is part of the conclusions in [12], which also discusses the changes in the shape of the isosceles trapezoid as $m_{2}$ changes. We note that reference [18] examines configurations considered in item (b) of proposition 5.4, as well as some symmetrical configurations with $m_{4}\neq m_{5}$ (none of which turns out to be central).
@@ 6. Generalized Dziobek Equations
The number of parameters defining $S$ is $c+1\choose 2$}, which, according to formula (A.2.4) in appendix A, is precisely the number of constraints that must be imposed on the mutual distances of a configuration of $n$ points to guarantee it has a given codimension $c>0$. This coincidence is not fortuitous since, by theorem A.2.2, the vanishing of $c+1\choose 2$} Cayley-Menger determinants provide the necessary constraints. Dziobek’s approach via critical point theory then introduces $c+1\choose 2$} Lagrange multipliers. Dziobek also verified that the derivative of a Cayley-Menger determinant with respect to a squared mutual distance can be expressed simply as a constant times the product of a pair of oriented volumes of suitable subconfigurations [16, 21]. Aside from determining the exact value of the constant, this fact can be readily inferred from theorem 6.1. Our proof of theorem 6.1 was inspired by the elegant proof of proposition 8 in [4].